Top Company holds 90 percent of Bottom Company’s common stock. In the current year, Top reports sales of $870,000 and cost of goods sold of $652,500. For this same period, Bottom has sales of $327,000 and cost of goods sold of $196,200. During the current year, Bottom sold merchandise to Top for $118,000. The parent still possesses 25 percent of this inventory at the current year-end. Bottom had established the transfer price based on its normal gross profit rate. What are the consolidated sales and cost of goods sold?
Answer #1

Solution: $1,079,000 and $742,500

Working Notes:

(1-(196,200/ 327,000) * 118,000 = $47,200

47,200*25% = $11,800

870000 + 327000 -118000 = $1,079,000

652,500 + 196,200 +11,800 - 118,000 = $742,500